Table 1 Pretreatment Covariates.

Covariate  Abbreviation Number of Levels

(if categorical)

Age  Age
Weight Index  WtI
Performance rating  PF  4
Cardiovascular disease history  HX  2
Systolic Blood pressure  SBP
Diastolic blood pressure  DBP
Electrocardiogram code EKG 7
Serum haemoglobin  HG
Size of primary tumour  SZ
Index of tumour stage and histolic grade SG
Serum prostatic acid phosphatase  AP
Bone metastases  BM  2

Table 2 Models and Fits.

Model Variable Groups #Parameters Log-likelihood Δ log L/
+11386.265 Δ# P
[LInd] - 55 0.000

58.77

[BPr] {SBP,DBP} 57 117.542

5.31

[3,2] {BM,WtI,HG},{SBP,DBP} 63 149.419

1.65

[5] {BM,WtI,HG,SBP,DBP} 75 169.163

1.31

[9]  Complement of {PF,HX,EKG}  127  237.092

Table 3a Clusters and outcomes for the untreated patients.

Outcome

Prostate Cardio Other
Patient Group  Alive  Death Death Death
Cluster 1 Stage 3  39  18  37  33
Stage 4  3  4  3  3
Cluster 2 Stage 3  1  4  2  3
Stage 4  14  49  18  6

Table 3b Clusters and outcomes for the treated patients.

Outcome

Prostate Cardio Other
Patient Group  Alive  Death Death Death
Cluster 1 Stage 3  50  3  52  20
Stage 4  4  0  1  3
Cluster 2 Stage 3  1  6  3  1
Stage 4  25  37  22  10